OpenStudy (anonymous):

What is the sum of the interior angles of a convex nonagon? A. 1620° B. 1800° C. 1260° D. 3240°

OpenStudy (alexandervonhumboldt2):

use formula (n-2)*180 where n represents number of sides: in your case n=9 (9-2)*180=7*180=1260